You have been given the following information about the production of Usher Co., and are asked to provide the plant manager with information for a meeting with the vice president of operations. Standard Cost Card Direct materials (5 pounds at $4 per pound) $20.00 Direct labor (0.80 hours at $10) 8.00 Variable overhead (0.80 hours at $3 per hour) 2.40 Fixed overhead (0.80 hours at $7 per hour) 5.60 $36.00 The following is a variance report for the most recent period of operations. Variances Costs Total Standard Cost Price Quantity Direct materials $410,000 $2,095 F $9,000 U Direct labor 164,000 3,906 U 22,000 U How many units were produced during the period? Number of units How many pounds of raw materials were purchased and used during the period? Raw material pounds What was the actual cost per pound of raw materials? (Round answer to 2 decimal places, e.g. 1.25.) Actual cost per pound of raw materials $ /lb How many actual direct labor hours were worked during the period? Actual direct labor hours hours What was the actual rate paid per direct labor hour? (Round answer to 2 decimal places, e.g. 1.25.) Actual rate paid per direct labor hour $ /hr
Answer #1
Computation of Actual Cost per Pound of Raw Material
Actual Quantity of Raw Material Used = (Direct Material Standard Cost + Direct Material Quantity Variance)/Stds. Price per Pound
=($410000+9000)/$4= 104750 Pound
Actual Cost per Pound =   (Direct materials Std. Cost +Direct materials quantity Variance -Direct materials price variance)/    actual Quantity of raw materials
=($410000+9000-2095)/ 104750= $3.98 per Pound
Hence , Acutal Cost of Material paid per pound = $3.98/Pound
Computation of Actual Rate per Direct Labour Hour
Actual Hour Used = (Direct Labour Standard Cost + Direct Labour Quantity Variance)/Stds. Rate per Hour
=($164000+22000)/$10= 18600Hour
Actual Rate per Hour =   (DL Std. Cost +DL quantity Variance + DL price variance)/    actual Hour Used
=($164000+22000+3906)/18600= $10.21 per Hour
Hence , Actual Rate per Hour paid = $10.21/Hour